OpenStudy (anonymous):

2/3 * 1/5? I have tofind the sum and simplify if possible help?

OpenStudy (anonymous):

@Pine We have to find the least common denominator, or the smallest number 3 and 5 can go into.

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

And please no direct answers.

OpenStudy (anonymous):

I wasn't asking for the answer I was asking for how to do it..

OpenStudy (anonymous):

@Pine What is the smallest number that 5 and 3 can both go into?

OpenStudy (anonymous):

Ok so the lcm would be 15?

OpenStudy (anonymous):

Yes.

OpenStudy (catlover5925):

take 3 * 5 to get the common denominator

OpenStudy (anonymous):

So you would multiply the denominators to a number that would equal 15. So what would you multiply 3 to to get 15?

OpenStudy (anonymous):

@Pine

OpenStudy (catlover5925):



OpenStudy (anonymous):

I would multiply 5 to get to 15

OpenStudy (anonymous):

Yes, but you would have to multiply 5 to the numerator also, to keep it balanced.

OpenStudy (anonymous):

So 13/15?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

Wait wait. I'm so confused

OpenStudy (anonymous):

@Pine Listen to me please. Since you multiply 5 to 3, you also have to multiply 5 to 2 to keep the fraction balanced.

OpenStudy (anonymous):

5 x 2 = ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

So 2/3 * 1/5= 10/15* 3/15??

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Hold on. So we got our first fraction, which is 10/15. Now we need the second fraction.

OpenStudy (anonymous):

Wait, yes you are correct.

OpenStudy (anonymous):

Can you correct my mistake there then?

OpenStudy (anonymous):

Oh crap.

OpenStudy (anonymous):

1/5x2/3 can be changed to 10/15x3/15 so 13/30

OpenStudy (anonymous):

I can't believe I forgot.

OpenStudy (anonymous):

You have to multiply straight across.

OpenStudy (anonymous):

(FaceDesk)

OpenStudy (anonymous):

So 10/15 * 3/15 so I got it right the very first time.?

OpenStudy (anonymous):

\(\Huge\frac{2}{3}\Huge{x}\frac{1}{5}\)

OpenStudy (anonymous):

Actually, you just multiply straight across.

OpenStudy (anonymous):

So It's 2 x 1 for the numerator, and 3 x 5 for the denominator.

OpenStudy (anonymous):

30/15????????? Or am I incredibly un-coherent?

OpenStudy (anonymous):

I said multiply straight across...

OpenStudy (anonymous):

My gosh my brain is dead

OpenStudy (anonymous):

2 x 1. that's the numerator. 3 x 5. That's the denominator.

OpenStudy (anonymous):

2/15

OpenStudy (anonymous):

Yup. :/

OpenStudy (anonymous):

Sorry, just a little confused because it went to 10/15 to 2/15

OpenStudy (anonymous):

Thanks I guess

OpenStudy (anonymous):

ya thats wrong

OpenStudy (anonymous):

What..

OpenStudy (anonymous):

I'm going to ask this again..

OpenStudy (anonymous):

3/25 is the answer i did reaserch

OpenStudy (anonymous):

That's wrong

OpenStudy (anonymous):

haha its funny because its not i used a calculator which is a little bit more3 accurate then your human calculator